1. Which element has the electron configuration:

1s2 2s2 2p6 3s2 3p6 4s2 3d10 4p6 5s2 4d5 ?

a. Br

b. Mn

c. Nb

d. Tc

2. What is the value of 27oC on the Fahrenheit scale?

Group of answer choices

a. -6.8

b. 106

c. 300

d. 81

e. none of these

3.

Which of the following atoms has the largest radius?

Group of answer choices

a. S

b. Se

c. Br

d. O
Answer #1

Qn1.

According to the electron configuration the element has 43 electrons. So the atomic number of the element is also 43. This is the atomic number of Technetium.

Technetium of Tc has this electron configuration.

Qn2.

We know, C /5=( F -32)/9

where, C = Temperature in celcius

F= Temperature in Fahrenheit

So, F -32 = (27/5)×9

F = 80.6 or ~81

So the answer is 81 .

Qn3.

Se atom has the largest radius.

Selenium has the largest atomic radius.